# Multimodal Graph Benchmark

### v1

@@ Appendix B Link Prediction Evaluation
Appendix C Experimental details
@@
Table 5: Best performing hyperparameters under Optuna search on Amazon-Sports, Amazon-Cloth and Goodreads-LP. *OOM = out of GPU memory.
Encoder Amazon-Sports Amazon-Cloth Goodreads-LP
Image Text LR # Layers LR # Layers LR # Layers
MMGCN CLIP 0.002595055216 1 0.00120638677883 2 OOM OOM
ViT T5 0.00372938084 1 0.00153969042325 2 OOM OOM
ImageBind 0.0008025598304 1 0.00051054426806 1 OOM OOM
Dinov2 T5 0.001184411766 2 0.00157538063161 2 OOM OOM
MGAT CLIP 0.001706100258 1 0.000109845035707 3 OOM OOM
ViT T5 0.001673818621 3 0.000242709240618 3 OOM OOM
ImageBind 0.000221326243 2 0.0001492753647 2 OOM OOM
Dinov2 T5 0.0001216520274 2 0.00007883247879 3 OOM OOM
GCN CLIP 0.00199258857 1 0.001250668422 1 0.0004599971993 1
ViT T5 0.0004748389636 1 0.0003867480145 1 0.0002695086535 1
ImageBind 0.0002173991668 1 0.00025841621630 1 0.0003440800285 1
Dinov2 T5 0.0003867480145 1 0.00050429857558 1 0.0005790300826 2
SAGE CLIP 0.000888801759 3 0.0006242779054 3 0.001115782634 3
ViT T5 0.0009069080338 3 0.0009159332036 3 0.001115782634 3
ImageBind 0.0007536062461 3 0.001115782634 3 0.0004026026641 3
Dinov2 T5 0.0009159332036 3 0.0009159332036 3 0.001115782634 3
BUDDY CLIP 0.0001996185501 1 0.0001073125661 2 0.0003131657128 1
ViT T5 0.0001782348253 1 0.000104734288 1 0.000101783244 1
ImageBind 0.0001082982111 3 0.0001048823969 3 0.001856957299 1
Dinov2 T5 0.0002695086535 1 0.0001385185507 1 0.0001996185501 1
MLP CLIP 0.0005982646116 1 0.0004259767004 1 0.0002317306399 3
ViT T5 0.00027681242 3 0.0003867480145 1 0.0001156508232 3
ImageBind 0.0001511847491 2 0.0001730467441 2 0.000101783244 3
Dinov2 T5 0.0002317306399 3 0.0003867480145 1 0.0001156508232 3
Table 6: Best performing hyperparameters under Optuna search on Ele-fashion and Goodreads-NC. *OOM = out of GPU memory.
Encoder Ele-fashion Goodreads-NC
Image Text LR # Layers LR # Layers
MMGCN CLIP 0.01617331603 1 OOM OOM
ViT T5 0.0001215688562 2 OOM OOM
ImageBind 0.01590607512 2 OOM OOM
Dinov2 T5 0.01233773648 1 OOM OOM
MGAT CLIP 0.0003211703387 1 OOM OOM
ViT T5 0.0004549756184 2 OOM OOM
ImageBind 0.003739550104 2 OOM OOM
Dinov2 T5 0.0002344834419 1 OOM OOM
GCN CLIP 0.002747501509 2 0.0004599971993 1
ViT T5 0.0007123650663 2 0.0002695086535 1
ImageBind 0.0007123650663 2 0.0003440800285 1
Dinov2 T5 0.0007123650663 2 0.0005790300826 2
SAGE CLIP 0.003235718844 3 0.001115782634 3
ViT T5 0.001523236953 2 0.001115782634 3
ImageBind 0.00174700527 2 0.0004026026641 3
Dinov2 T5 0.00134217448 2 0.001115782634 3
MLP CLIP 0.004407317752 3 0.00174700527 3
ViT T5 0.001342211088 2 0.0004414532347 3
ImageBind 0.00278539483 2 0.0001409760217 3
Dinov2 T5 0.001287758261 3 0.0004611586099 3
C.1 Hyperparameter Tuning
@@ C.2 Optimization
The best-performing hyperparameters for each combination of feature encoders and GNNs are presented in Table 5 and 6. These hyperparameters are selected based on the highest mean performance across the 3 runs.
C.3 Reproducibility
We make the source code for MM-Graph publicly available on GitHub, along with detailed documentation and instructions for running the experiments. The datasets used in MM-Graph are also made accessible through the Hugging Face Datasets platform, allowing researchers to easily access and use the data for their own experiments.
By providing a comprehensive experimental setup, detailed hyperparameter configurations, and open-source code and datasets, we aim to foster reproducibility and encourage further research on multimodal graph learning using the MM-Graph benchmark.
Appendix D Dataset
In this section, we provide details on the release and maintenance of the Multimodal Graph Benchmark (MM-Graph) datasets, the file structure of the datasets, and instructions on how to use them effectively.
D.1 Dataset Release
To facilitate easy access and widespread use of the MM-Graph datasets, we have created a dedicated webpage at https://mm-graph-benchmark.github.io/. The webpage provides a direct link to the datasets and showcases representative examples from each dataset, allowing researchers to quickly grasp the nature and diversity of the data.
The MM-Graph datasets are hosted on the Hugging Face Datasets platform at https://huggingface.co/datasets/mm-graph-org/mm-graph. Hugging Face Datasets is a popular and reliable platform for hosting and sharing datasets in the machine learning community. By leveraging this platform, we ensure the long-term maintenance and preservation of the MM-Graph datasets. We plan to periodically update the repository to address any issues or incorporate feedback from the community.
The MM-Graph datasets are released under the Apache 2.0 license, which allows for open access and permissive use of the data for research purposes. This licensing choice aligns with our goal of fostering collaborative research and encouraging the development of advanced multimodal graph learning algorithms.
To complement the datasets, we also provide a standardized benchmark, evaluator, and various GNN backbones and feature encoders on our GitHub repository at https://github.com/mm-graph-benchmark/mm-graph-benchmark. The repository includes detailed documentation and example usage to facilitate the adoption and extension of the MM-Graph benchmark.
D.2 Dataset File Structure
The MM-Graph directory contains five separate folders, each corresponding to a specific dataset: amazon-sports, amazon-cloth, goodreads-lp, goodreads-nc, and ele-fashion. Each folder contains the necessary files for the respective dataset. For the three link prediction datasets (amazon-sports, amazon-cloth, and goodreads-lp), each folder contains the following files:
  • lp-edge-split.pt: This file contains the graph structures, along with the train, validation, and test edges, as well as their corresponding negative pairs.
  • clip_feat.pt, t5vit_feat.pt, imagebind_feat.pt, t5dino_feat.pt: These files contain the node features obtained from various feature encoders, as discussed in Section 4.2 of the paper.
For the two node classification datasets (goodreads-nc and ele-fashion), each folder contains the following files:
  • edges.pt: This file contains the graph structures.
  • split.pt: This file indicates the train, validation, and test nodes.
  • label-w-missing.pt: This file stores the node labels. Nodes with missing labels are assigned to a dummy class and are not used as train, validation, or test nodes.
  • clip_feat.pt, t5vit_feat.pt, imagebind_feat.pt, t5dino_feat.pt: These files contain the node features obtained from various feature encoders, as discussed in Section 4.2 of the paper.
D.3 Dataset Usage
To load and use the MM-Graph datasets, we provide a user-friendly Python package.
For detailed examples and usage instructions, please refer to the GitHub repository. The repository provides comprehensive documentation, including installation instructions, dataset descriptions, and example scripts for training and evaluating models on the MM-Graph datasets.
By providing a standardized and easy-to-use interface for accessing the datasets, along with detailed documentation and examples, we aim to facilitate the adoption of the MM-Graph benchmark and enable researchers to focus on developing and evaluating their multimodal graph learning models.

### v2

@@ Appendix B Link Prediction Evaluation
Appendix C Other Candidate Datasets
We also provide a list of candidate datasets and the reasons why they are unsuitable for the our benchmark here.
Ogbn-Arxiv: Raw paper IDs are not provided and cannot be retrieved by backtracking, making it impossible to crawl images from Arxiv.
Yelp: Only reviews of restaurants have textual and visual information. However, the graph structure of user-friend-with-user exists, and typically, a user’s friends do not have any reviews on any restaurant. Additionally, Yelp’s license prevents repurposing the datasets.
Tiktok: This dataset does not provide any raw text or image features.
Appendix D Experimental details
@@
D.1 Hyperparameter Tuning
@@ D.2 Optimization
D.3 Additional Results
SOTA methods applied to text-only data. We compare ImageBind+SAGE with UniGraph, one of the SOTA methods on text-attributed graphs, and results shown in Tab. 8 are from  [14]. It demonstrates that multimodal inputs consistently outperform text-only inputs.
Table 8: Node classification results on Goodreads-NC and Ele-Fashion.
Dataset UniGraph ImageBind+SAGE
Amazon-Sports 27.11 $\pm$ 0.10 34.32 $\pm$ 0.11
Amazon-Cloth 18.01 $\pm$ 0.03 25.20 $\pm$ 0.09
Goodreads-LP 22.31 $\pm$ 0.05 34.61 $\pm$ 0.43
Goodreads-NC 78.14 $\pm$ 0.11 80.39 $\pm$ 0.21
Ele-fashion 81.05 $\pm$ 0.08 87.71 $\pm$ 0.13
F1 score
Additionally, we report F1-score on the two node classification datasets. The results are shown below.
Table 9: Additional results on Macro-F1 score
Image Encoder Text Encoder Ele-fashion Goodreads-NC
MMGCN CLIP CLIP 75.07 $\pm$0.45 81.02 $\pm$0.32
T5 ViT 63.98 $\pm$1.18 79.45 $\pm$0.49
ImageBind ImageBind 76.18 $\pm$0.72 78.41 $\pm$0.45
T5 DINOv2 73.62 $\pm$0.04 80.13 $\pm$0.13
GCN CLIP CLIP 79.85 $\pm$0.04 74.36 $\pm$0.01
T5 ViT 79.69 $\pm$0.08 76.27 $\pm$0.03
ImageBind ImageBind 80.47 $\pm$0.12 72.84 $\pm$0.10
T5 DINOv2 79.39 $\pm$0.09 81.72 $\pm$0.08
SAGE CLIP CLIP 87.08 $\pm$0.01 83.30 $\pm$0.04
T5 ViT 85.81 $\pm$0.04 82.94 $\pm$0.08
ImageBind ImageBind 87.72 $\pm$0.08 80.26 $\pm$0.04
T5 DINOv2 85.45 $\pm$0.12 82.88 $\pm$0.02